OpenStudy (anonymous):

- 1/3 + w = 5 3/8 what is w? @iGreen

OpenStudy (igreen):

Well we have to add 1/3 to both sides. What's \(\dfrac{1}{3} + 5\dfrac{3}{8}\)

OpenStudy (anonymous):

not so sure @iGreen

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

PLEASE @iGreen

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

PLEASE SOMEBODY!!!!!!!!

OpenStudy (igreen):

First convert 5 3/8 into an improper fraction. Multiply the denominator to the whole number, then add it to the numerator. \(5 \dfrac{3}{8} \rightarrow \dfrac{8 \times 5 + 3}{8} \rightarrow \dfrac{43}{8}\) So now we have: \(\dfrac{1}{3} + \dfrac{43}{8}\) Now we have to find a common denominator between 3 and 8..

OpenStudy (igreen):

Do you know what that will be?

OpenStudy (anonymous):

24 @iGreen

OpenStudy (anonymous):

Is that correct @iGreen

OpenStudy (igreen):

Yep, you got it. Now we change both fractions to denominators of 24. \(\dfrac{1}{3} \rightarrow \dfrac{1 \times 8}{3 \times 8} \rightarrow \dfrac{8}{24}\) \(\dfrac{43}{8} \rightarrow \dfrac{43 \times 3}{8 \times 3} \rightarrow \dfrac{129}{24}\) Now can you add \(\dfrac{8}{24} + \dfrac{129}{24}\)?

OpenStudy (anonymous):

137/24

OpenStudy (igreen):

Yep, you got it. Now we have to switch it back to a mixed number.

OpenStudy (anonymous):

5

OpenStudy (igreen):

Yes, that would be the whole number part, but we still need the fraction.

OpenStudy (anonymous):

5 17/24

OpenStudy (igreen):

Yep, you got it! Our final answer will be \(5 \dfrac{17}{24}\).

OpenStudy (anonymous):

THANK YOU SOOOO MUCH! I have like 2 more can u help me

OpenStudy (igreen):

Yeah, I might be able to..just tag me.

OpenStudy (anonymous):

THNX! @iGreen

OpenStudy (texaschic101):

or this way.. -1/3 + w = 5 3/8 -1/3 + w = 43/8 --- multiply everything by 24 -8 + 24w = 129 24w = 129 + 8 24w = 137 w = 137/24 w = 5 17/24

OpenStudy (texaschic101):

if you multiply by the common denominator, it will get rid of the fractions

OpenStudy (igreen):

Ooo, that looks much easier..

OpenStudy (texaschic101):

either way you should arrive at the same answer

OpenStudy (anonymous):

Thank you too @texaschic101